# From Mono- to Hexa-Interstitials: Computational Insights into Carbon Defects in Diamond

### v1

@@ III.4 Electronic Properties
The mono-interstitial adopts a singlet ground state with a wavefunction of $D_{2d}$ symmetry. Its $a_{1}$ state is fully occupied and lies within the valence band (VB). In this case, the presence of two dangling bonds introduced by the $p$-orbital of the two $sp^{2}$ coordinated neighboring carbon atoms gives rise to a pair of degenerate $e$ levels in each spin channel within the band gap. In an ideal symmetric singlet, one electron occupies each spin channel, distributed equally between the two components of the $e$ manifold. This behavior is notably reproduced by the PBE functional: the two Kohn-Sham (KS) levels remain strictly degenerate, and each carries an occupation of $\approx 0.5$ per spin channel, corresponding to a half-filled $e$ doublet with the correct singlet occupation. In contrast, the HSE06 hybrid functional lifts this degeneracy. Although the two states still transform as an $E$ representation, one component ($e_{x}$ in spin-up and $e_{y}$ in spin-down) is lowered in energy and becomes fully occupied, while the complementary component ($e_{y}$ in spin-up and $e_{x}$ in spin-down) is shifted upward by roughly 3 eV and remains empty. Thus, HSE06 drives the system from the symmetric half-half occupation characteristic of PBE to an integer-occupation solution in which one member of the $e$ pair is filled, and the other is depopulated. Thus, the KS level splitting is not a consequence of a Jahn–Teller distortion nor of any underlying lattice-symmetry breaking; rather, it reflects the hybrid functional’s intrinsic tendency to penalize fractional occupations, together with the increased level of separation introduced by Fock-exchange.
@@
Based on our KS level analysis, none of the lowest-energy carbon self-interstitial configurations exhibits the electronic structure required to account for the TR12 center. The TR12 center has an experimental ZPL of $\sim$2.63 eV, implying that any viable defect candidate must possess two well-isolated defect states whose KS energy separation exceeds this value. No such pair of in-gap states is found for the neutral self-interstitial configurations considered here. Although their formation energies suggest that these lowest-energy defects can exist in multiple charge states, which is interesting in its own right, we do not pursue them further here.
@@ III.5 Vibrational Properties
where $v_{n,\alpha}$ denotes the eigenvector component of defect-center atom $n$ along direction $\alpha$, which is a real number at the $\Gamma$-point. Partial intensities are then obtained by scaling the total mode activities by the corresponding $\rho$ values. Defect-center atoms are defined as those that are strongly displaced from their ideal crystallographic sites; their numbers are 6, 6, 8, 12, 14, and 18 atoms for the $\mathrm{C}^{a}_{\mathrm{1i}}$$\mathrm{C}^{a}_{\mathrm{6i}}$ complexes, respectively.
@@
The IR spectra clearly exhibit many distinct peaks, indicating that IR spectroscopy is highly sensitive to self-interstitials and can reliably discriminate between different defect types. Whereas a pristine diamond is IR-inactive at the zone center because inversion symmetry forbids dipole-active phonons, interstitial defects, especially those introducing local $sp^{2}$ bonding, break this symmetry and activate vibrational modes with nonzero dipole moments. This symmetry breaking enables a range of IR-active features. Many of the stiffer, defect-induced bonds give rise to high-frequency IR modes with well-separated and clearly distinguishable peaks. There are only four peaks at low frequency and localized on the defect; two at 470 and 518 cm-1 in $\mathrm{C}^{a}_{\mathrm{1i}}$, and one in both $\mathrm{C}^{a}_{\mathrm{5i}}$ and $\mathrm{C}^{a}_{\mathrm{6i}}$ at 465 and 474 cm-1, respectively. All high-energy IR-active modes are fully localized on the bonds associated with the defect centers, as illustrated in Fig. 5(c).
@@ IV Conclusions
We have carried out a systematic investigation of carbon self-interstitial complexes in diamond, spanning from mono- to hexa-interstitials. To overcome the severe configurational complexity that emerges for larger clusters, we employed interatomic potential-assisted searches, followed by full relaxation and energetics within DFT. Accurate hybrid functional calculations were used to ensure reliable defect stability ordering and defect-level alignments. This multiscale protocol enabled us to produce all established low-energy configurations for small interstitials and to identify several previously unreported tri-, penta-, and hexa-interstitial geometries.
@@
Our results confirm that the neutral $\langle 001\rangle$-oriented split interstitial is the only stable mono-interstitial configuration in the electronic ground state, while the commonly reported lower-symmetry alternative may arise from the limitations of semi-local functionals. For di-interstitials, we recover the known stability hierarchy, with the $\pi$-bonded complex as the ground state. For higher-order complexes, we find a clear thermodynamic drive toward aggregation: the formation energy per interstitial decreases with cluster size, showing a pronounced drop from isolated and di-interstitials to compact tri- and tetra-interstitials. The tetra-interstitial platelet is predicted to be a particularly stable motif, consistent with its proposed role as a building block for extended $\{100\}$ planar platelets. As additional carbon atoms are incorporated, the bond-length window remains narrow, yet the spread in bond strengths increases, reflecting the coexistence of very stiff defect-core bonds and weaker defect–lattice linkages.
@@
The electronic structures reveal two distinct classes of self-interstitial complexes: mono-, di-, penta-, and hexa-interstitials can host localized in-gap states and multiple charge states, associated with dangling bonds or partly reconstructed motifs. In contrast, all low-energy tri- and tetra-interstitials are electronically silent: they introduce no gap states and remain stable only in the neutral charge state. Crucially, none of the lowest-energy neutral self-interstitial complexes exhibits the pair of well-isolated in-gap states required to explain the TR12 center with its $\sim$2.63 eV ZPL. This rules out neutral self-interstitial clusters as viable TR12 candidates and suggests that alternative intrinsic or impurity-assisted defects must be responsible for the experimentally observed center.
@@
Vibrational spectroscopy provides complementary fingerprints. In Raman spectra, the lattice-phonon region remains broadly similar across defects, aside from systematic blue shifts originating from internal strain induced by fixed pristine lattice vectors. Defect-center vibrations generate additional Raman peaks at high frequency, but their non-resonant first-order activities are negligible. By contrast, IR spectra show rich, defect-specific signatures. Local symmetry breaking, often accompanied by $sp^{2}$-like rehybridization, activates dipole-allowed modes that are absent in pristine diamond. Short, stiff C-C bonds within the defect cores yield well-separated high-frequency IR peaks (1375-1925 cm-1), and mode-projection analysis confirms that these features are strongly localized on defect-center bonds. These results establish IR spectroscopy as a particularly powerful probe for discriminating among self-interstitial complexes.
@@
Overall, this work delivers a unified and functionally robust picture of the stability, electronic activity, and vibrational signatures of carbon self-interstitial aggregates in diamond.

### v2

@@ III.4 Electronic Structure
The mono-interstitial adopts a singlet ground state with a wave function of $D_{2d}$ symmetry. Its $a_{1}$ state is fully occupied and lies within the valence band (VB). In this case, the presence of two dangling bonds introduced by the $p$-orbital of the two $sp^{2}$ coordinated neighboring carbon atoms gives rise to a pair of degenerate $e$ levels in each spin channel within the band gap.
@@
In an ideal symmetric singlet, one electron occupies each spin channel, distributed equally between the two components of the $e$ manifold. This behavior is notably reproduced by the PBE functional: the two KS levels remain strictly degenerate, and each carries an occupation of $\approx 0.5$ per spin channel, corresponding to a half-filled $e$ doublet with the correct singlet occupation. In contrast, the HSE06 hybrid functional lifts this degeneracy. Although the two states still transform as an $E$ representation, one component ($e_{x}$ in spin-up and $e_{y}$ in spin-down) is lowered in energy and becomes fully occupied, while the complementary component ($e_{y}$ in spin-up and $e_{x}$ in spin-down) is shifted upward by roughly 3 eV and remains empty. Thus, HSE06 drives the system from the symmetric half-half occupation characteristic of PBE to an integer-occupation solution in which one member of the $e$ pair is filled, and the other is depopulated. We emphasize that the KS level splitting is not a consequence of a Jahn–Teller distortion nor of any underlying lattice-symmetry breaking; rather, it reflects the hybrid functional’s intrinsic tendency to penalize fractional occupations, together with the increased level of separation introduced by Fock-exchange. This electron configuration instead corresponds to one of the Slater determinants of the $^{1}B_{1}$ multiplet [31]. Promoting a spin-down electron from $e_{y}$ to $e_{x}$ yields a mixture of the $^{1}A_{1}$ and $^{1}B_{2}$ multiplet states, which are optically inactive in the high-symmetry configuration but may become dynamically activated through coupling to symmetry-distorting phonons [31]. Possible correlations of this defect with near-infrared and ultraviolet optical centers are discussed in Ref. [31]. Since the main focus of this work is to understand the properties of carbon self-interstitial clusters, we continue our investigation with the di-interstitial complexes.
@@
All the most stable clusters that introduce mid-gap states and undergo electron excitation to the conduction band lead to bound-exciton excitations. These transitions typically occur in the ultraviolet regime; therefore, such defects are not considered in the present study. Based on our KS level analysis, none of the lowest-energy carbon self-interstitial cluster configurations exhibits the electronic structure required to account for the TR12 center. The TR12 center has an experimental ZPL of $\sim$2.63 eV, implying that any viable defect candidate must possess two well-isolated defect states whose KS energy separation exceeds this value. No such pair of in-gap states is found for the neutral self-interstitial cluster configurations considered here. Although their formation energies suggest that these lowest-energy defects can exist in multiple charge states, which is interesting in its own right, we do not pursue them further here.
@@ III.5 Vibrational Properties
where $v_{n,\alpha}$ denotes the eigenvector component of the defect-center atom $n$ along the direction $\alpha$, which is a real number at the $\Gamma$-point. Partial intensities are then obtained by scaling the total mode activities by the corresponding $\rho$ values. Defect-center atoms are defined as those that are strongly displaced from their ideal crystallographic sites; their numbers are 6, 6, 8, 12, 14, and 18 atoms for the $\mathrm{C}^{a}_{\mathrm{1i}}$$\mathrm{C}^{a}_{\mathrm{6i}}$ complexes, respectively.
@@
The IR spectra clearly exhibit many distinct peaks, indicating that IR spectroscopy is highly sensitive to self-interstitials and can reliably discriminate between different defect types. Whereas a pristine diamond is IR-inactive at the zone center because inversion symmetry forbids dipole-active phonons, interstitial defects, especially those introducing local $sp^{2}$ bonding, break this symmetry and activate vibrational modes with nonzero dipole moments. This symmetry breaking enables a range of IR-active features. Many of the stiffer, defect-induced bonds give rise to high-frequency IR modes with well-separated and clearly distinguishable peaks. There are only four peaks at low frequency, localized on the defect; two at 470 and 518 cm-1 in $\mathrm{C}^{a}_{\mathrm{1i}}$, and one in both $\mathrm{C}^{a}_{\mathrm{5i}}$ and $\mathrm{C}^{a}_{\mathrm{6i}}$ at 465 and 474 cm-1, respectively. All high-energy IR-active modes are fully localized on the bonds associated with the defect centers, as illustrated in Fig. 5(c).
III.6 Identification of 3H center and its relation with TR12 center
Figure 6: (a) The Kohn–Sham levels of the open-shell singlet state of the neutral $\mathrm{C}^{b}_{\mathrm{2i}}$ defect, represented by spin-up ($\uparrow$) and spin-down ($\downarrow$) channels. Representation of the defect-state orbitals for (b) an open-shell (OS) singlet broken-symmetry wave function and (c) a closed-shell (CS) symmetric wave function treatment. The light cyan and yellow lobes exhibit negative and positive isovalue, with an absolute value of $5\times 10^{-7}$ Å-3.
The experimental detection of the R1 EPR center, associated with the di-interstitial $\mathrm{C}^{c}_{\mathrm{2i}}$, indicates that metastable configurations can likewise appear as local minima on the potential energy surface [54, 64]. To this end, we also consider configurations that can form and persist under kinetic control. Among these, the $\mathrm{C}^{b}_{\mathrm{2i}}$ defect (the Humble di-interstitial) is particularly compelling because it introduces both filled and empty levels in the fundamental band gap of diamond.
As mentioned, Goss et al. proposed that the 3H optical center corresponds to $\mathrm{C}^{b}_{\mathrm{2i}}$ in the positive charge state based on the computed local vibration modes that were compared to experiments [31, 65]. Our formation-energy diagram showed that under optical excitation, the positively charged defect would ionize to the neutral charge state, so it would not remain stable at the relevant excitation energy. Therefore, this attribution appears to be incorrect. On the other hand, the neutral charge state cannot be disregarded as a viable candidate which is photostable under blue illumination.
The neutral $\mathrm{C}^{b}_{\mathrm{2i}}$ defect forms a two-state system, as shown in Fig. 6(a). It exhibits a broken-symmetry singlet ground state, as seen from orbitals visualization in Fig. 6(b). In order to test this assertion, we performed symmetric closed-shell (CS) calculations (Fig. 6(c)), which show that its energy is 0.56 eV higher than that obtained from our OS singlet calculations. In a broken-symmetry system, the wave functions spontaneously break the space-group symmetry, and the KS orbitals are not necessarily eigenstates of the symmetry operations of the defective supercell [66]. Such a dilemma can be alleviated by adopting a multiplet-state representation  [67, 68]. Considering the $a_{2}$ and $b_{1}$ defect orbitals in $C_{2v}$ symmetry, four states are obtained:
$\Psi\!\left({}^{1}A_{1}\right)$ $=\big[a_{2}(1)a_{2}(2)\big]\big[\uparrow\downarrow-\downarrow\uparrow\big],$ (10a)
$\Psi\!\left({}^{1}A_{1}\right)$ $=\big[b_{1}(1)b_{1}(2)\big]\big[\uparrow\downarrow-\downarrow\uparrow\big],$ (10b)
$\Psi\!\left({}^{1}B_{2}\right)$ $=\big[a_{2}(1)b_{1}(2)+b_{1}(1)a_{2}(2)\big]\big[\uparrow\downarrow-\downarrow\uparrow\big],$ (10c)
$\Psi\!\left({}^{3}B_{2}\right)$ $=\big[a_{2}(1)b_{1}(2)-b_{1}(1)a_{2}(2)\big]\left\{\begin{array}[]{c}\big[\uparrow\uparrow\big]\\ \big[\uparrow\downarrow+\downarrow\uparrow\big]\\ \big[\downarrow\downarrow\big]\end{array}\right\}.$ (10g)
Equation 10a represents the CS singlet state. Equation 10b corresponds to a double-electron excitation configuration, which lies $2.58$ eV higher in energy than the ground state. However, the relevant OS singlet ground state arises when one electron occupies the spin-up channel of the $a_{2}$ orbital and the other occupies the spin-down channel of the $b_{1}$ orbital (or vice versa), corresponding to the $a_{2}\otimes b_{1}={}^{1}B_{2}$ configuration, as given in Eq. 10c. This state is described by a broken-symmetry wave function that reproduces the ground-state density of the multiplet state. The corresponding triplet state, $^{3}B_{2}$, given by Eq. 10g, lies 0.13 eV higher in energy.
The relevant excitation of the 3H center transforms the ground state [Eq. (10c)] into the excited state [Eq. (10b)] via a double-excitation process, yielding a ZPL of 2.58 eV, in good agreement with the experimental value of 2.46 eV [65]. A single-electron excitation cannot describe this transition, as it does not recover the multideterminant, spin-adapted singlet configuration required for an accurate many-body description. In principle, the two CS $^{1}A_{1}$ states should be treated as linear combinations (i.e., mixed). However, because the calculated energy separation between the configurations in Eqs. (10c) and (10b) is $\approx 2.0$ eV, we expect the mixing to be weak and therefore regard our approach as a reasonable approximation.
[…]
Figure 7: The PL spectrum of the neutral $\mathrm{C}^{b}_{\mathrm{2i}}$ defect compared with the experimental data from Ref. [65]. A rigid shift of $0.12$ eV is applied to the calculated spectrum to align the zero-phonon line with experiment. The inset highlights the evolution of the PL spectrum as a function of isotope mass. The experiments were carried out for a diamond sample with a 50%-50% mixture of 13C and 12C isotopes (red spectrum).
The PL spectrum of the 3H optical center was also measured in a specially prepared diamond sample containing a 50%–50% mixture of 13C and 12C isotopes [65]. This isotopic disorder gives rise to additional features in the PL spectrum; most notably, a pronounced triple peak emerges in the phonon sideband (PSB) around 2.25 eV, which has been associated with local vibrational modes of $\sim$230 meV. Isotope-induced shifts and splittings in the PSB provide a sensitive means to validate microscopic models of optical centers. In addition to the energies of the local modes, their relative intensities in the PL spectrum are also informative, and are accessible within the Huang–Rhys framework used here.
Isotope enrichment systematically renormalizes vibrational frequencies: increasing the average atomic mass lowers phonon frequencies and compresses the PSB toward the ZPL. To reproduce the experimentally observed triple-peak structure in the 2.24–2.26 eV range, we generated 50 random isotope configurations and report the ensemble-averaged lineshape. Throughout, we assume that isotope substitution leaves the interatomic force constants essentially unchanged; accordingly, only the atomic masses were modified when constructing the dynamical matrix (see, e.g., Ref. [69]). The dynamical matrix was then solved separately for each configuration.
Mass substitution results in an approximately 10 meV shift of the PSB. The characteristic triple-peak structure is most pronounced at a 50% 13C concentration, yielding excellent agreement with experiment. This agreement provides additional support for assigning the 3H center to the neutral $\mathrm{C}^{b}_{\mathrm{2i}}$ defect. Finally, the experimentally observed peak at 2.15 eV, which is absent from our calculations, likely originates from residual laser light or a different color center rather than a 3H-related emission feature.
We note that the 3H center is particularly interesting because it features a singlet ground state and an experimentally confirmed photoactive singlet excited state [65], together with a low-lying triplet manifold. This level structure could be exploited for qubit operation, where the metastable triplet state may serve as an optically addressable qubit. However, the large energy separation between the optically excited singlet and the metastable triplet is expected to suppress intersystem crossing (ISC), potentially limiting spin polarization and readout. One possible route to enhance ISC is a two-color excitation scheme, in which a second laser resonantly drives vibronic transitions to access spin–orbit-coupled channels more efficiently, thereby enabling qubit operation.
Figure 8: (a) The Kohn–Sham energy levels of the open-shell singlet state of the neutral $\mathrm{C}^{b}_{\mathrm{6i}}$ defect, shown separately for the spin-up ($\uparrow$) and spin-down ($\downarrow$) channels. (b) Schematic illustration of the electron excitation mechanisms investigated in the search for the TR12 PL signal. (c) The PL spectra of the neutral $\mathrm{C}^{b}_{\mathrm{6i}}$ defect compared with the experimental data taken from Ref. [18]. For better comparison with the experiment, rigid energy shifts of $0.10$ eV and 0.42-0.42 eV were applied to the calculated spectra to align the ZPL with the experimental value. Our calculated ZPL for the occupation numbers of 0.1 and 0.5 are 2.53 eV and 3.06 eV, respectively.
By establishing this identification and directly overlaying the experimentally measured PL spectra of the 3H and TR12 centers, several notable similarities become apparent. Beyond the proximity of their ZPLs—2.46 eV for 3H and 2.63 eV for TR12—their phonon sidebands (PSBs) exhibit multiple common spectral features [18, 65]. Furthermore, annealing studies show opposing trends in their intensities: the TR12 signal generally increases with annealing, whereas the 3H signal decreases [12, 70]. This anticorrelation suggests that TR12 may form as a secondary product associated with the transformation of the 3H center. Motivated by this possibility, we systematically screened our defect dataset for more complex configurations that intrinsically incorporate the Humble structural motif (i.e., $\mathrm{C}^{b}_{\mathrm{2i}}$). This search ultimately highlighted the $\mathrm{C}^{b}_{\mathrm{6i}}$ defect as a prime candidate.
The $\mathrm{C}^{b}_{\mathrm{6i}}$ defect is composed of the $\mathrm{C}^{b}_{\mathrm{2i}}$ and $\mathrm{C}^{a}_{\mathrm{4i}}$. While the isolated $\mathrm{C}^{a}_{\mathrm{4i}}$ defect does not introduce localized states within the band gap and instead produces a deep valence-band state, it plays a crucial role when incorporated into the $\mathrm{C}^{b}_{\mathrm{6i}}$ configuration. Here we label the character of the electronic states as those for the isolated Humble $\mathrm{C}^{b}_{\mathrm{2i}}$ defect with $C_{2v}$ symmetry for the sake of simplicity; the symmetry of the composed defect is rather $C_{1h}$. As shown in Fig. 8(a), the $\mathrm{C}^{a}_{\mathrm{4i}}$-related state, labeled $a_{1}$, shifts toward the valence-band maximum and thereby may participate in the electronic excitation process. When considering the $a_{2}$ and $b_{1}$ states, this configuration exhibits an electronic structure similar to that of the $\mathrm{C}^{b}_{\mathrm{2i}}$ defect, where the singlet ground state is described by a broken-symmetry solution. The singlet–triplet energy difference is 60 meV (see Table 2).
For two electrons occupying a four-state system, the many-body states can be constructed in the same manner as for the 3H configuration discussed above. The broken-symmetry singlet ground state, $a_{2}\otimes b_{1}=B_{2}$, is followed by the closed-shell $a_{2}\otimes a_{2}=A_{1}$ state, which lies 0.73 eV higher in energy. The $^{3}B_{2}$ triplet state is found at 0.06 eV above the broken-symmetry singlet state. The $b_{1}\otimes b_{1}=A_{1}$ configuration arises from a double-electron excitation and lies 2.46 eV above the ground state. However, incorporating the two extra states originating from $\mathrm{C}^{a}_{\mathrm{4i}}$ leads to a six-state system with 15 many-body configurations (six singlets and three triplets with nine components), reflecting more complexity.
To identify the relevant excitation process, we use the experimental ZPL (2.63 eV) and HRF (2.30) of the TR12 center as reference criteria. As illustrated schematically (Fig. 8), one of the lowest optical transitions, denoted ex1, corresponds to a double-electron excitation from the $a_{2}$ to the $b_{1}$ level, yielding a computed ZPL of 2.46 eV and an HRF of 3.22. Another excitation, ex2, involves a single-electron promotion from the $a_{1}$ to the $b_{1}$ state, for which we obtain a ZPL of 3.52 eV and an HRF of 0.84. Comparison of these two candidate excitations suggests that the system behaves as a strongly interacting two-electron two-hole manifold, in which the actual optical transition may arise from a mixture of the ex1 and ex2 configurations rather than from a pure single excitation. To probe this possibility, we introduced fractional occupations of the two excited configurations as a proxy for a mixed many-body state (see ex3 in Fig. 8(b)). Specifically, we considered fractional occupations of $(0.1/0.9)$ and $(0.5/0.5)$ for the two excitation components.
Figure 8(c) shows the PL lineshapes. For the $(0.5/0.5)$ occupation, we obtain a ZPL of 3.06 eV and an HRF of 1.46. The $(0.1/0.9)$ occupation yields a ZPL of 2.53 eV with an HRF of 2.32, which is closer to the experimental values. However, the corresponding PSB structure does not fully reproduce the experimental spectral features. We attribute this discrepancy either to limitations of the computational methodology employed in this work or to finite-size effects arising from the relatively small supercell used for the six-interstitial defect.
We propose that $\mathrm{C}^{b}_{\mathrm{6i}}$ may represent a potential candidate for the TR12 center; however, a definitive assignment requires higher-level excited-state methods capable of simultaneously capturing electron–electron correlation and vibronic coupling in the excited state. Such an investigation using multireference many-body approaches is beyond the scope of the present work. Nevertheless, based on the energetic ordering, ZPL alignment, and qualitative agreement of the PL spectral features, we conclude that the $\mathrm{C}^{b}_{\mathrm{6i}}$ defect remains a viable candidate for the TR12 center.
@@ IV Conclusions
We have carried out a systematic investigation of carbon self-interstitial complexes in diamond, spanning from mono- to hexa-interstitial defects. To overcome the severe configurational complexity that emerges for larger clusters, we employed interatomic potential-assisted structure searches, followed by full relaxation and energetics within DFT. Accurate hybrid functional calculations were used to ensure reliable defect stability ordering and defect-level alignments. This multiscale protocol enabled us to produce all established low-energy configurations for small interstitials and to identify several previously unreported tri-, penta-, and hexa-interstitial geometries.
@@
Our results confirm that the neutral $\langle 001\rangle$-oriented split interstitial is lowest-energy mono-interstitial configuration, while the commonly reported lower-symmetry alternative arises from the limitations of semi-local functionals. For di-interstitials, we recover the known stability hierarchy, with the $\pi$-bonded complex as the ground state. For higher-order complexes, we find a clear thermodynamic drive toward aggregation: the formation energy per interstitial decreases with cluster size, showing a pronounced drop from isolated and di-interstitials to compact tri- and tetra-interstitials. The tetra-interstitial platelet is predicted to be a particularly stable motif, consistent with its proposed role as a building block for extended $\{100\}$ planar platelets. As additional carbon atoms are incorporated, the bond-length window remains narrow, yet the spread in bond strengths increases, reflecting the coexistence of very stiff defect-core bonds and weaker defect–lattice linkages.
@@
The electronic structures reveal two distinct classes of self-interstitial complexes. Mono-, di-, penta-, and hexa-interstitials can host localized in-gap states and multiple charge states, associated with dangling bonds or partly reconstructed motifs. In contrast, all low-energy tri- and tetra-interstitials are electronically silent; they introduce no gap states and remain stable only in the neutral charge state, highlighting their relevance as optically and electrically silent defects in irradiated and annealed diamond. Crucially, none of the lowest-energy neutral self-interstitial complexes exhibits the pair of well-isolated in-gap states required to account for the TR12 center with its $\sim$2.63 eV ZPL. This observation rules out the most stable self-interstitial aggregates as microscopic models of TR12. Motivated by this discrepancy, we extended our analysis to metastable configurations that may form and persist under kinetic control. Within this broader configurational landscape, our investigation of metastable configurations identifies the 3H defect center as a neutral di-interstitial defect. Building on this identification, we further propose that the TR12 center may originate from a 3H-containing defect linked to a metastable hexa-interstitial configuration.
@@
Vibrational spectroscopy provides complementary fingerprints of defects and might be the only experimental method for exploring electronically inactive defects. For the most stable complexes we analyze the Raman and infrared signatures associated with these defects. In Raman spectra, the lattice-phonon region remains broadly similar across defects, aside from systematic blue shifts originating from internal strain induced by fixed pristine lattice vectors. Defect-center vibrations introduce additional Raman peaks at high frequency, but their non-resonant first-order activities are negligible. By contrast, IR spectra show rich, defect-specific signatures. Local symmetry breaking, often accompanied by $sp^{2}$-like rehybridization, activates dipole-allowed modes that are absent in pristine diamond. Short, stiff C-C bonds within the defect cores yield well-separated high-frequency IR peaks (1375-1925 cm-1), and mode-projection analysis confirms that these features are strongly localized on defect-center bonds. These results establish IR spectroscopy as a particularly powerful probe for discriminating among self-interstitial complexes.
@@
Overall, this work delivers a unified and functionally robust picture of the stability, electronic activity, and vibrational signatures of carbon self-interstitial aggregates in diamond , and establishes a consistent framework for linking first-principles defect models with experimental identification.